OpenStudy (anonymous):

A rectangular box is twice as long as it is wide. The height of the box is 3 feet less than the width. If the box is x feet wide, what polynomial represents its volume in cubic feet? A.4x-3 B. 2x^3+6x^2 C.2x^2-6x D.2x^3-6x^2

OpenStudy (anonymous):

V=L*W*H W=x H=x-3 L=2W